OpenStudy (anonymous):

6x + y = 13 y - x = -8 Solve By Elimination.

OpenStudy (anonymous):

6x + y = 13 y - x = -8 multiply by -1 -y+x = 8 add it to the other equation 7x = 21 x = 3 y-3=-8 y=-5

OpenStudy (anonymous):

medal?

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

ur welcome